# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v1

@@ Appendix A. Supplementary material
Description of the experimental setup; plots of the mean-square displacement of Li+, TFSI-, and CO2 within the electrolyte obtained from our MD simulations; plots showing the variation of positive electrode over-potential during the discharge process at various current densities as a function of the state of discharge; a table of parameters used in the finite element analysis (FEA); a table of boundary conditions used in FEA; a table of initial values for independent and dependent variables used in FEA; and plots illustrating changes in porosity and volume fraction of deposited species in the cathode throughout the discharge process at different current densities and discharge states.

### v5

@@ Appendix A. Supplementary material
Description of the experimental setup; a table of parameters used in the finite element analysis (FEA); a table of boundary conditions used in FEA; a table of initial values for independent and dependent variables used in FEA; plots showing the voltage-capacity curve during Li-CO2 battery discharge at various rates ranging from 0.2 mA/cm2 to 0.9 mA/cm2; plots showing the effect of initial values of relevant variables on the voltage-capacity curve during Li-CO2 battery discharge at a rate of 0.4 mA/cm2; plots illustrating changes in porosity and volume fraction of deposited species in the cathode throughout the discharge process at different current densities and discharge states; and plots showing the change in CO2 concentration within the cathode.